OpenStudy (anonymous):

How many ways are there to choose a committee of 5 people from a group of 15 people? How do you solve this what is the equation?

terenzreignz (terenzreignz):

15C5 Like, fifteen, choose five. ^_^ You know the formula for nCr?

OpenStudy (anonymous):

no I dont

terenzreignz (terenzreignz):

You know what "factorial" means?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Okay, good. Then the formula for nCr is as follows: \[\Large _nC_r = \frac{n!}{r!\cdot (n-r)!}\] Just plug in :)

OpenStudy (anonymous):

is n 5 and 4 15?

terenzreignz (terenzreignz):

What do you think? ;) in general, n should be the larger number.

terenzreignz (terenzreignz):

Did you get your answer?

OpenStudy (anonymous):

is 3,003

terenzreignz (terenzreignz):

That's right ^^ Remember, that's what you do whenever you're asked how many ways you can pick r distinct members from a set of n. nCr. Good job :)

OpenStudy (anonymous):

why can't you use nPr

terenzreignz (terenzreignz):

Good question. nPr \[\Large _nP_r = \frac{n!}{(n-r)!}\] is used for when ORDER matters. For instance, if the question were how many ways we can ARRANGE five members from a group of 15 people in a ROW, then the answer would involve nPr instead.

OpenStudy (anonymous):

thanks